# PrivTuner with Homomorphic Encryption and LoRA: A P3EFT Scheme for Privacy-Preserving Parameter-Efficient Fine-Tuning of AI Foundation Models

### v1

@@ Abstract
AI foundation models have recently demonstrated impressive capabilities across a wide range of tasks. Fine-tuning (FT) is a method of customizing a pre-trained AI foundation model by further training it on a smaller, targeted dataset. In this paper, we initiate the study of the Privacy-Preserving Parameter-Efficient FT (P3EFT) framework, which can be viewed as the intersection of Parameter-Efficient FT (PEFT) and Privacy-Preserving FT (PPFT). PEFT modifies only a small subset of the model’s parameters to achieve FT (i.e., adapting a pre-trained model to a specific dataset), while PPFT uses privacy-preserving technologies to protect the confidentiality of the model during the FT process. There have been many studies on PEFT or PPFT, but very few on their fusion, which motivates our work on P3EFT to achieve both parameter efficiency and model privacy. To exemplify our P3EFT, we present the PrivTuner scheme, which incorporates Fully Homomorphic Encryption (FHE) enabled privacy protection into LoRA (short for “Low-Rank Adapter”). Intuitively speaking, PrivTuner allows the model owner and the external data owners to collaboratively implement PEFT with encrypted data. After describing PrivTuner in detail, we further investigate its energy consumption and privacy protection. Then, we consider a PrivTuner system over wireless communications and formulate a joint optimization problem to adaptively minimize energy while maximizing privacy protection, with the optimization variables including FDMA bandwidth allocation, wireless transmission power, computational resource allocation, and privacy protection. A resource allocation algorithm is devised to solve the problem. Experiments demonstrate that our algorithm can significantly reduce energy consumption while adapting to different privacy requirements.
@@ I Introduction
Privacy-Preserving Parameter-Efficient FT: P3EFT studied in this paper. We initiate the study of Privacy-Preserving Parameter-Efficient Fine Tuning (P3EFT) and propose a Fully Homomorphic Encryption (FHE) based scheme called PrivTuner for P3EFT (we will elaborate on FHE in later paragraphs). P3EFT can be understood as the intersection of PEFT and PPFT. Particularly, our proposed PrivTuner fuses LoRA and FHE technologies to provide efficient and privacy-enhanced fine-tuning services to external entities. We will detail the system model of PrivTuner in Sec. IV.
@@
  • We initiate the study of the P3EFT framework to bridge the gap between PEFT and PPFT.
  • To show a specific instance of P3EFT, we design a novel scheme named PrivTuner, where the model owner and external devices collaboratively implement secure and efficient fine-tuning tasks. Specifically, PrivTuner is supported by LoRA and FHE, which could significantly reduce the training burden and ensure both model and data security.
@@
  • A resource allocation (RA) algorithm employing both branch and bound (B&B) and fractional programming techniques is devised to solve the formulated problem effectively. Time complexity, solution quality and convergence analysis are also provided. Experimental results validate the effectiveness and superiority of our algorithm.
@@ II-A Low-Rank Adapters (LoRA) for AI Foundation Models
The motivation of LoRA [10] is that the updates to pre-trained model weights are likely to have a low intrinsic rank during adaption. Hence, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the update is constrained by a low-rank decomposition: $W_{0}+\triangle W=W_{0}+A^{1}A^{2}$, where $A^{1}\in\mathbb{R}^{d\times r},A^{2}\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min\{d,k\}$. During training, $W_{0}$ is frozen and not updated, while $A^{1}$ and $A^{2}$ contain trainable parameters. When given an input $x$ to the model, the forward pass yields:
$y=W_{0}x+\triangle x=W_{0}x+A^{1}A^{2}x.$ (1)
@@ II-B Fully Homomorphic Encryption (FHE)
  • $\mathrm{Dec}(\mathrm{sk},\widetilde{c})\rightarrow{m}$: On secret key $\mathrm{sk}$ and a ciphertext $\widetilde{c}$, perform encryption to obtain the plaintext message $m$.
  • $\mathrm{Eval}(\mathrm{pk},\widetilde{c}_{1},\widetilde{c}_{2},\mathrm{func})\rightarrow\hat{c}$: On public key $\mathrm{pk}$, two ciphertexts $\widetilde{c}_{1},\widetilde{c}_{2}$ encrypted from $m_{1},m_{2}$, and a linear function $\mathrm{func}$, outputs a new ciphertext $\widetilde{c}$ such that it is the same as $\mathrm{Enc}(\mathrm{pk},\mathrm{func}(m_{1},m_{2}))$.
A more detailed discussion of CKKS can be found in [24].
@@ III-A Discussing Related Studies Based on Figure 1(a)
After proposing the term P3EFT, we found a recent unpublished work [25] which also coins P3EFT independently. The study [25] incorporates split learning [26] into PEFT to achieve only heuristic privacy without even defining what protection is being offered (in fact, split learning is shown to be inherently insecure [27]). In addition, Bu et al. [28] incorporate the formal notion of differential privacy (DP) into PEFT to achieve P3EFT, but the name P3EFT is not mentioned and DP is orthogonal to the privacy protection mechanism of HE used in our paper. Also, DP inherently introduces noise or randomness to the model and may degrade the model’s accuracy, while HE generally does not hurt accuracy [29]. Finally, the idea of P3EFT (but not the name) is also briefly discussed in [30] as a potential research direction.
@@ III-B Discussing Related Studies Based on Figure 1(b)
As noted in the subsection above, only a few studies [25, 28, 30] touch upon P3EFT, and none of them is about RA, so there is no very related work on RA of PrivTuner or other P3EFT instances. Hence, except [25, 28, 30], there is no other work closely related to PrivTuner. Yet, since PrivTuner is an important contribution of our paper, we present some work on “FHE + Deep Learning (DL)” from the perspective of different kinds of DL models. The main arithmetic components in neural networks are matrix addition and multiplication, which are linear and easily supported by FHE. The DL models to be discussed include CNNs, RNNs, and transformers.
@@ IV System Model
This section presents the system model, beginning with the threat model and an overview in Sec. IV-A. Then, we elaborate on the proposed scheme of PrivTuner in Sec. IV-B. Fig. 2 presents the overall flows of PrivTuner.
@@ IV-A Threat Model and Overview
Consider a system comprising a model owner’s server and a set of external mobile devices, denoted as $\mathcal{N}:=\{1,\ldots,N\}$. The server hosts a pre-trained foundation model serving as a Machine Learning as a Service (MLaaS) platform (e.g., ChatGPT operated by OpenAI). Each mobile device $n\in\mathcal{N}$ aims to fine-tune the foundation model using its own private data to enhance personalized performance. Operating under an honest but curious security model [47], our PrivTuner allows devices to transmit FHE-encrypted data instead of the raw one to the server. Then, the server creates an adapter for each device via LoRA, and the inference results (also encrypted) are returned to the devices. Upon reception, devices decrypt the inference results and calculate the loss with the corresponding labels. The loss information is then sent back to the server for gradient calculation and adapter updating.
IV-B PrivTuner
@@
3for each mobile device $n\in\mathcal{N}$ in parallel do
@@
5      $\mathrm{KeyGen}(\lambda_{n},q_{n})$rightarrow $\mathrm{pk}_{n},\mathrm{sk}_{n}$;
6      $\mathrm{Enc}(\mathrm{pk}_{n},X_{n}^{\textnormal{ft}})$rightarrow $\widetilde{X}_{n}^{\textnormal{ft}}$;
@@ 7      Transmit $\widetilde{X}_{n}^{\textnormal{ft}}$ and $\mathrm{pk}_{n}$ to the server;
11Generate adapters: $\mathrm{LoRA}_{n}({W}_{0})\rightarrow\mathcal{A}_{n}:=\{A_{n}^{1},A_{n}^{2}\},~{}\forall n\in\mathcal{N}$;
12Step 3: Prediction.
13Output predictions: $\widetilde{Y}_{n}^{\textnormal{p}}=\mathrm{Eval}\big{(}{\mathrm{pk}}_{n},({W}_{0},\mathcal{A}_{n}),\widetilde{X}_{n}^{\textnormal{ft}},f^{p}\big{)},~{}\forall n\in\mathcal{N}$, where $f^{p}$ is the prediction function defined in (6);
14Send result $\widetilde{Y}_{n}^{\textnormal{p}}$ back to each device $n$;
15Step 4: Decryption and Loss Computation.
16for each mobile device $n\in\mathcal{N}$ in parallel do
17
18      $\mathrm{Dec}(\mathrm{sk}_{n},\widetilde{Y}_{n}^{\textnormal{p}})$rightarrow $Y_{n}^{\textnormal{p}}$;
19      Compute the loss: $\mathcal{L}_{n}(Y_{n}^{\textnormal{p}},Y_{n}^{\textnormal{ft}})$rightarrow $L_{n}$;
@@ 20      Transmit $L_{n}$ to the server;
22
23Step 5: Update of Adapters.
Update each adapter $\mathcal{A}_{n}$ with $L_{n}$ and obtain $\mathcal{A}^{\textnormal{ft}}_{n}:=\{A_{n}^{1,\textnormal{ft}},A_{n}^{2,\textnormal{ft}}\}$.
Algorithm 1 Steps of PrivTuner.
@@
where $\lambda_{n}$ is the polynomial degree and $q_{n}$ is the coefficient modulus. With the generated public key $\mathrm{pk}_{n}$ from (2), now device $n$ could encrypt its fine-tuning data $X_{n}$ as follows:
@@
Adapter Generation. Upon receiving $\widetilde{X}_{n}$ and $\mathrm{pk}_{n}$, the server could apply various PEFT methods to the pre-trained foundation model $W_{0}$. Here we employ LoRA [10] (explained in Sec. II-A) to generate an adapter $\mathcal{A}_{n}=\{A_{n}^{1},A_{n}^{2}\}$ for each device $n$ as described below:
@@
where $\mathrm{LoRA}_{n}(\cdot)$ denotes the low-rank adaption operation (we have the subscript “$n$” just for generality, and the experiments use the same for all $n$). Since the size of $\mathcal{A}_{n}$ is much smaller than that of $W_{0}$, this approach dramatically reduces the number of trainable parameters.
Prediction. Then, the server could perform predictions on the encrypted data $\widetilde{X}_{n}^{\textnormal{ft}}$ with $(W_{0},\mathcal{A}_{n})$ and $\mathrm{pk}_{n}$, generating the corresponding predictions $\widetilde{Y}_{n}^{\textnormal{p}}$ (also encrypted):
@@
$f^{p}({W}_{0},A_{n}^{2},\widetilde{X}_{n}^{\textnormal{ft}})=W_{0}\widetilde{X}_{n}^{\textnormal{ft}}+A_{n}^{1}A_{n}^{2}\widetilde{X}_{n}^{\textnormal{ft}},~{}\forall n\in\mathcal{N}.$ (6)
Since the fine-tuning labels $\bm{Y}^{\textnormal{ft}}$ are kept by mobile devices, the server sends $\widetilde{Y}_{n}^{\textnormal{p}}$ back to each device $n$ for loss computation. It is worth noting that using prediction $\widetilde{Y}_{n}^{\textnormal{p}}$ to reason about the parameter matrix $W_{0}$ is computationally infeasible, especially for devices with limited computing resources. Thus, this will not compromise security.
@@
Update of Adapters. Given that the loss $L_{n}$ and weight matrices $W_{0},\mathcal{A}_{n}=\{A_{n}^{1},A_{n}^{2}\}$ are in plaintext, the server can compute the gradient and update $\mathcal{A}_{n}$ into fine-tuned $\mathcal{A}_{n}^{\textnormal{ft}}$ following a standard fine-tuning process. For an arbitrary input $X_{n}^{\textnormal{input}}$, now the server could provide the prediction service as:

### v3

@@ Abstract
AI foundation models have recently demonstrated impressive capabilities across a wide range of tasks. Fine-tuning (FT) is a method of customizing a pre-trained AI foundation model by further training it on a smaller, targeted dataset. In this paper, we initiate the study of the Privacy-Preserving Parameter-Efficient FT (P3EFT) framework, which can be viewed as the intersection of Parameter-Efficient FT (PEFT) and Privacy-Preserving FT (PPFT). PEFT modifies only a small subset of the model’s parameters to achieve FT (i.e., adapting a pre-trained model to a specific dataset), while PPFT uses privacy-preserving technologies to protect the confidentiality of the model during the FT process. There have been many studies on PEFT or PPFT, but very few on their fusion, which motivates our work on P3EFT to achieve both parameter efficiency and model privacy. To exemplify our P3EFT, we present the PrivTuner scheme, which incorporates Fully Homomorphic Encryption (FHE) enabled privacy protection into LoRA (short for “Low-Rank Adapter”), a popular PEFT solution published in ICLR 2021 [1]. Intuitively speaking, PrivTuner allows the model owner and the external data owners to collaboratively implement PEFT with encrypted data. After describing PrivTuner in detail, we further investigate its energy consumption and privacy protection. Then, we consider a PrivTuner system over wireless communications and formulate a joint optimization problem to adaptively minimize energy while maximizing privacy protection, with the optimization variables including FDMA bandwidth allocation, wireless transmission power, computational resource allocation, and privacy protection. A resource allocation algorithm is devised to solve the problem. Experiments demonstrate that our algorithm can significantly reduce energy consumption while adapting to different privacy requirements.
@@ I Introduction
Privacy-Preserving Parameter-Efficient FT: P3EFT studied in this paper. We initiate the study of Privacy-Preserving Parameter-Efficient Fine Tuning (P3EFT). P3EFT can be understood as the intersection of PEFT and PPFT. To demonstrate the feasibility of P3EFT, we propose a specific instantiation, PrivTuner, which leverages Fully Homomorphic Encryption (FHE) and LoRA for privacy-preserving and computation-efficient fine-tuning. By integrating FHE, PrivTuner enables computations to be performed directly on encrypted data, preventing data exposure while maintaining the efficiency of LoRA-based adaptation. This design allows external devices to fine-tune models collaboratively with minimal trust assumptions, making it suitable for scenarios such as federated learning, cloud-based AI services, and edge intelligence. We will detail the system model of PrivTuner in Section IV.
Resource Allocation in PrivTuner. Resource Allocation (RA) plays a pivotal role in the wireless communication network where resources (e.g., transmission power and bandwidth) are often limited. In our paper, we consider PrivTuner to operate in a wireless communication network comprising one server and multiple clients, and RA could effectively manage different resources to improve the overall network performance. By incorporating RA, we can formulate a customized optimization problem for the PrivTuner network that aligns with specific resource budgets and performance demands (e.g., energy consumption, latency and privacy) and then design an efficient algorithm to solve it. The problem formulation and RA algorithm design are discussed in Sections VI and VII.
@@
  • We introduce the P3EFT framework, which bridges the gap between PEFT and PPFT by integrating their advantages to enable both efficient and privacy-preserving fine-tuning. To validate this framework, we propose PrivTuner as a concrete instance of P3EFT. PrivTuner combines LoRA with FHE to provide secure and efficient collaborative fine-tuning between the model owner and external devices, reducing the computational burden while ensuring data and model privacy.
@@
  • A resource allocation algorithm employing both branch and bound (B&B) and fractional programming techniques is devised to solve the formulated problem effectively. Time complexity, solution quality and convergence analysis are also provided. Experimental results validate the effectiveness and superiority of our algorithm.
@@ II-A Low-Rank Adapters (LoRA) for AI Foundation Models
The motivation of LoRA [1] is that the updates to pre-trained model weights are likely to have a low intrinsic rank during adaption. Hence, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the update is constrained by a low-rank decomposition: $W_{0}+\Delta W=W_{0}+A^{1}A^{2}$, where $A^{1}\in\mathbb{R}^{d\times r},A^{2}\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min\{d,k\}$. During training, $W_{0}$ is frozen and not updated, while $A^{1}$ and $A^{2}$ contain trainable parameters. When given an input $x$ to the model, the forward pass yields:
$y=W_{0}x+\Delta Wx=W_{0}x+A^{1}A^{2}x.$ (1)
@@ II-B Fully Homomorphic Encryption (FHE)
  • $\mathrm{Dec}(\mathrm{sk},\widetilde{c})\rightarrow{m}$: On secret key $\mathrm{sk}$ and a ciphertext $\widetilde{c}$, perform decryption to obtain the plaintext message $m$.
  • $\mathrm{Eval}(\mathrm{pk},\widetilde{c}_{1},\widetilde{c}_{2},\mathrm{func})\rightarrow\widetilde{c}$: On public key $\mathrm{pk}$, two ciphertexts $\widetilde{c}_{1},\widetilde{c}_{2}$ encrypted from $m_{1},m_{2}$, and a linear function $\mathrm{func}$, outputs a new ciphertext $\widetilde{c}$ such that it is the same as $\mathrm{Enc}(\mathrm{pk},\mathrm{func}(m_{1},m_{2}))$.
A more detailed discussion of CKKS can be found in [24]. Specifically, CKKS is designed for approximate arithmetic on real and complex numbers, distinguishing it from integer-based schemes (e.g., BFV, BGV). In addition, CKKS effectively manages the noise growth through customized scaling and rescaling techniques, which could prevent the overflow issue. The above characteristics make CKKS well-suited for privacy-preserving deep learning tasks. Deep learning models often operate on real-valued numbers (e.g., floating-point numbers) rather than exact integers. These models are also robust to small numerical errors or approximations due to the reliance on optimization techniques and probabilistic learning. Consequently, CKKS is widely used in privacy-preserving applications across various AI models, including CNNs [25, 19, 20], RNNs [26, 27], and Transformers [28, 29].
@@ III-A Discussing Related Studies based on Figure 1(a)
After proposing the term P3EFT, we found a recent unpublished work [30] which also coins P3EFT independently. The study [30] incorporates split learning [31] into PEFT to achieve only heuristic privacy without even defining what protection is being offered (in fact, split learning is shown to be inherently insecure [32]). In addition, Bu et al. [33] incorporate the formal notion of differential privacy (DP) into PEFT to achieve P3EFT, but the name P3EFT is not mentioned and DP is orthogonal to the privacy protection mechanism of HE used in our paper. Also, DP inherently introduces noise or randomness to the model and may degrade the model accuracy, while Homomorphic Encryption (HE) generally does not hurt accuracy [34]. Finally, the idea of P3EFT (but not the name) is also briefly discussed in [35] as a potential research direction.
@@ III-B Discussing Related Studies based on Figure 1(b)
As noted in the subsection above, only a few studies [30, 33, 35] touch upon P3EFT, and none of them is about RA, so there is no very related work on RA of PrivTuner or other P3EFT instances. Hence, except [30, 33, 35], there is no other work closely related to PrivTuner. Yet, since PrivTuner is an important contribution of our paper, we present some work on “FHE + Deep Learning (DL)” from the perspective of different kinds of DL models. The DL models to be discussed include CNNs, RNNs, and transformers.
@@
Resource allocation has been widely studied in the context of wireless communications, where it is often employed to optimize the use of network resources under different constraints. Previous works have applied RA to federated learning [45, 46], semantic communication [47, 48] and unmanned aerial vehicle (UAV) frameworks [49, 50], showing its effectiveness in improving overall network performance under different application scenarios.
@@ IV System Model
This section presents the system model, beginning with the threat model and an overview in Section IV-A. Then, we elaborate on the proposed scheme of PrivTuner in Section IV-C. Fig. 2 presents the overall flows of PrivTuner.
@@ IV-A Threat Model and Overview
Consider a system comprising a model owner’s server and a set of external mobile devices, denoted as $\mathcal{N}\mathrel{\mathop{\ordinarycolon}}=\{1,\ldots,N\}$. The server hosts a pre-trained foundation model serving as a Machine Learning as a Service (MLaaS) platform (e.g., ChatGPT operated by OpenAI). Each mobile device $n\in\mathcal{N}$ aims to fine-tune the foundation model using its own private data to enhance personalized performance. Operating under an honest but curious security model [57], our PrivTuner allows devices to transmit FHE-encrypted data instead of the raw one to the server. Then, the server creates an adapter for each device via LoRA, and the inference results (also encrypted) are returned to the devices. Upon reception, devices decrypt the inference results and calculate the loss with the corresponding labels. The loss information is then sent back to the server for gradient calculation and adapter updating.
IV-B HE-friednly BERT-Tiny Model
In our work, we adopt the BERT-Tiny model, a small and efficient variant of the original BERT [58]. Following the approach proposed by Rovida et al. [29], we use the CKKS scheme to perform encrypted computations while maintaining an acceptable trade-off between computational overhead and model accuracy.
The employed BERT-Tiny model involves large matrix multiplications and several non-linear layers (e.g., Softmax, GeLU, and LayerNorm). In particular, those non-linear functions can not be implemented directly, as the current CKKS only support addition and multiplication operations. To efficiently handle matrix multiplications in ciphertext, the SIMD technique supported by CKKS could be leveraged, which allows parallel computations on encrypted vectors. For different non-linear functions involved, we follow Rovida et al. [29] and employ different methods (e.g., Maclaurin series and Chebyshev polynomial) to approximate them efficiently. To save space, more detailed information and the approximation performance for non-linear layers are reported in Appendix. A.
IV-C PrivTuner
@@
3 for each mobile device $n\in\mathcal{N}$ in parallel do
@@
5 $\mathrm{KeyGen}(\lambda_{n},q_{n})$ $\rightarrow$ $\mathrm{pk}_{n},\mathrm{sk}_{n}$;
6 $\mathrm{Enc}(\mathrm{pk}_{n},X_{n}^{\textnormal{ft}})$ $\rightarrow$ $\widetilde{X}_{n}^{\textnormal{ft}}$;
@@ 7 Transmit $\widetilde{X}_{n}^{\textnormal{ft}}$ and $\mathrm{pk}_{n}$ to the server;
11
12Generate adapters: $\mathrm{LoRA}_{n}({W}_{0})\rightarrow\mathcal{A}_{n}\mathrel{\mathop{\ordinarycolon}}=\{A_{n}^{1},A_{n}^{2}\},~\forall n\in\mathcal{N}$;
13Step 3: Prediction.
14Output predictions: $\widetilde{Y}_{n}^{\textnormal{p}}=\mathrm{Eval}\big({\mathrm{pk}}_{n},({W}_{0},\mathcal{A}_{n}),\widetilde{X}_{n}^{\textnormal{ft}},f^{p}\big),~\forall n\in\mathcal{N}$, where $f^{p}$ is the prediction function defined in (6);
15Send result $\widetilde{Y}_{n}^{\textnormal{p}}$ back to each device $n$;
16Step 4: Decryption and Loss Computation.
17for each mobile device $n\in\mathcal{N}$ in parallel do
18
19 $\mathrm{Dec}(\mathrm{sk}_{n},\widetilde{Y}_{n}^{\textnormal{p}})$ $\rightarrow$ $Y_{n}^{\textnormal{p}}$;
20 Compute the loss: $\mathcal{L}_{n}(Y_{n}^{\textnormal{p}},Y_{n}^{\textnormal{ft}})$ $\rightarrow$ $L_{n}$;
@@ 21 Transmit $L_{n}$ to the server;
23
24Step 5: Update of Adapters.
Update each adapter $\mathcal{A}_{n}$ with $L_{n}$ and obtain $\mathcal{A}^{\textnormal{ft}}_{n}\mathrel{\mathop{\ordinarycolon}}=\{A_{n}^{1,\textnormal{ft}},A_{n}^{2,\textnormal{ft}}\}$.
Algorithm 1 PrivTuner.
@@
where $\lambda_{n}$ is the polynomial degree and $q_{n}$ is the coefficient modulus. With the generated public key $\mathrm{pk}_{n}$ from (2), now device $n$ could encrypt its fine-tuning data $X_{n}^{\textnormal{ft}}$ as follows:
@@
Adapter Generation. Upon receiving $\widetilde{X}_{n}^{\textnormal{ft}}$ and $\mathrm{pk}_{n}$, the server could apply various PEFT methods to the pre-trained foundation model $W_{0}$. Here we employ LoRA [1] (explained in Section II-A) to generate an adapter $\mathcal{A}_{n}=\{A_{n}^{1},A_{n}^{2}\}$ for each device $n$ as described below:
@@
where $\mathrm{LoRA}_{n}(\cdot)$ denotes the low-rank adaption operation (we have the subscript “nn” just for generality, and the experiments use the same adapter for all $n$). Since the size of $\mathcal{A}_{n}$ is much smaller than that of $W_{0}$, this approach dramatically reduces the number of trainable parameters.
Prediction. Then, the server could perform predictions on the encrypted data $\widetilde{X}_{n}^{\textnormal{ft}}$ with $(W_{0},\mathcal{A}_{n})$ and $\mathrm{pk}_{n}$, generating the corresponding predictions $\widetilde{Y}_{n}^{\textnormal{p}}$ (also encrypted). As discussed in Section IV-B, we could follow Rovida et al. [29] to implement the prediction, and the process can be generalized as follows:
@@
$f^{p}({W}_{0},\mathcal{A}_{n},\widetilde{X}_{n}^{\textnormal{ft}})=W_{0}\widetilde{X}_{n}^{\textnormal{ft}}+A_{n}^{1}A_{n}^{2}\widetilde{X}_{n}^{\textnormal{ft}},~\forall n\in\mathcal{N}.$ (6)
where $W_{0}$ is the foundation model, $\mathcal{A}_{n}$ is the generated adapter, and $\widetilde{X}_{n}^{\textnormal{ft}}$ is the encrypted dataset for fine-tuning. Since the fine-tuning labels $\bm{Y}^{\textnormal{ft}}$ are kept by mobile devices, the server sends $\widetilde{Y}_{n}^{\textnormal{p}}$ back to each device $n$ for loss computation. It is worth noting that using prediction $\widetilde{Y}_{n}^{\textnormal{p}}$ to reason about the parameter matrix $W_{0}$ is computationally infeasible, especially for devices with limited computing resources. Thus, this will not compromise security.
@@
Updates. Given the loss $L_{n}$ and weight matrices $W_{0},\mathcal{A}_{n}=\{A_{n}^{1},A_{n}^{2}\}$, the server can compute the gradient and update $\mathcal{A}_{n}$ into fine-tuned $\mathcal{A}_{n}^{\textnormal{ft}}$ following a standard fine-tuning process [44]. For an arbitrary input $X_{n}^{\textnormal{input}}$, now the server could provide the prediction service as: